Friday, March 23, 2018 - 15:24

We are receiving reports of phone outages for multiple agencies. This is only affecting the phone capabilities. The internet appears to be unaffected at this time.

AZNET has already been notified and is working to resolve the issue. The issue is being tracked with remedy ticket 177260

Wednesday, February 21, 2018 - 11:56

AZNET Voice issues affecting long distance and 1-800 calls have been resolved successfully. 

If you encounter any further issues with the service, please contact the AZNET II Service Desk at (602) 364-4444, option 1.

The reference ticket number for this issue is Remedy 174202.

Wednesday, February 21, 2018 - 10:55

AZNET is currently experiencing issues with voice services related to long distance calls. This impacts multiple state agencies. Our AZNET teams are currently working to restore services as soon as possible. 

At this time we do not have an ETA. We will provide an update when available and apologize for this inconvenience.

AZNET II has opened a Customer Bridge number, you may call in at: 866-789-8818 PIN: 574769

Please reference master ticket: 174202.

Wednesday, February 21, 2018 - 10:53

We are still experiencing issues with users logging into Office 365 webmail. When attempting to log in, users receive a "Username or Password is incorrect" message. This is currently affecting multiple agencies. Our teams are currently working to restore services as soon as possible. 

At this time we do not have an ETA. We will provide an update when available and apologize for this inconvenience.

Please reference master ticket number: 36250 when calling the ADOA Service Desk.

Tuesday, February 20, 2018 - 09:39

We are currently experiencing issues with users logging into Office 365 webmail. When attempting to log in, users receive a "Username or Password is incorrect" message. This is currently affecting multiple agencies. Our teams are currently working to restore services as soon as possible. 

At this time we do not have an ETA. We will provide an update when available and apologize for this inconvenience.

Please reference master ticket number: 36250 when calling the ADOA Service Desk.

Thursday, February 15, 2018 - 15:20

Today Feb 15 starting at 3:10 PM we are getting reports that people cannot sign into HRIS/YES. HRIS Technical has been informed and are looking to provide a solution as soon as possible.

The issue can be tracked with ticket 35994

Wednesday, January 31, 2018 - 10:58

We are receiving reports that the Talent Acquisition pages are responding slowly. HRIS Technical has been informed and is currently working to resolve the latency.

The ticket number for this issue is 33888

[Resolved]

Wednesday, January 24, 2018 - 09:12

Around 8:55am, we started receiving reports issues with the Copper Wifi access not working. We have informed AZNET of the issue and they are currently working for a resolution.

This issue is being tracked with Remedy ticket 171896.

This has been resolved.

Sunday, January 21, 2018 - 10:11

Good morning, All -

We had a GO/NO GO Decision Call with all of the agency decision makers at 9:45 AM this morning, and I am delighted to report that we reached a unanimous GO decision from our agency partners!

With this in mind, we will now complete the final steps in fully migrating the ADOA Mainframe over to the multi-tenant zCloud environment in Boulder, Colorado, with a DR presence in Raleigh, North Carolina.

Thank you to all the many people who contributed to bringing this successful day to fruition.   It was truly a magnificent collaborative effort!

This will be the final posting for this morning, but a final email will be distributed to the project team containing all the pertinent information required going forward, including the number of the Conference Bridge that IBM will be opening as of 3:00 PM this afternoon, MST.

 

 

 

Sunday, January 21, 2018 - 09:02

Good morning - 

The time is 9:02 AM, and at the present time, we have had reports of completed successful testing from the following agencies:

DES, DPS, DOC, ASU (one issue with check writing remains, which is being looked at by the ASU network technician), and AHCCCS.

ADOT is waiting to hear on AMVA testing.   One issue for DOR was just resolved, and production jobs are now running at DOR, so they are close to completing.

We are now looking ahead to the Final GO/NO GO Decision Call this morning at 9:45 AM.

The next update will be posted immediately after this call to notify everyone of the final GO/NO GO decision!

 

 

Sunday, January 21, 2018 - 08:02

Good morning - 

At this hour, all agencies are busy with the user verification testing window, which will continue until 9:45 AM, when we will be making our final GO/NO GO decision call.

The IBM Operations team is resolving agency issues discovered in testing as they arise, and so far all is going well.

DES has already reported that their testing is done successfully, and has already provided their "GO" decision, but DES will be on the 9:45 AM call, as well, to make it official!

The next update will be posted at 9:00 AM.

 

Sunday, January 21, 2018 - 07:21

Good morning - 

Apologies for the lateness of this report - there has been a flurry of activity with the LPARs in the past hour!

So far, AHCCCS, DOR, DOC, ASU, DPS and ADOT have been given the go to start testing, all between 6:50 AM and 7:10 AM.

We are still awaiting word from DES.

The next update will be posted at 8:00 AM.

 

Sunday, January 21, 2018 - 06:00

Good morning - 

The time now is 6:00 AM.

The IODEF file copy issue mentioned in the 5:00 AM update has been resolved.

At the present time, the IBM Operations Team is in the process of bringing up the LPARs in Boulder.

We are currently, approximately 10 minutes ahead of schedule.

Agency representatives from AHCCCS and ADOT have arrived at the State Data Center.

We are currently still looking to begin the agency user verification window on schedule at 6:30 AM.

The next update will be posted at 7:00 AM.

 

Sunday, January 21, 2018 - 05:04

Good morning -

The time is now 5:04 AM.

We are currently slightly ahead of schedule.  All Phoenix LPARs were brought down at 4:17 AM.

The network has been swung over to Boulder, and the LPARs are in the process of being brought up there.

All VTAM work on the LPARs has been completed.

There is currently an issue with the copying of the IODEF file, that is being troubleshot by Shawn Beardsley and Tom Shaffer of IBM.

At this time, the Agency Conference Bridge has been opened.   All who would care to join can do so by dialing 1-877-820-7831, and adding the Participant Passcode of 410516#.

Our next update will be posted at 6:00 AM. 

Sunday, January 21, 2018 - 03:50

Good morning, All -

The time is now 3:52 AM on Sunday, January 21, 2018.   

At exactly 4:00 AM, we will officially begin our scheduled outage for the Final Cutover of the ADOA Mainframe to the multi-tenant zCloud environment in Boulder, Colorado.

The LPARs will begin to be brought down from Phoenix, conceivably for the final time, and brought back up in Boulder, with a user verification testing period that is scheduled to run from 6:30 AM until 9:45 AM.

At 9:45 AM, we will have a conference call with all of the agency stakeholders (decision makers) to obtain a FINAL GO/NO GO DECISION.   If we are a "GO," the data will continue its final replication to Boulder and the mainframe system will remain in Boulder from this time forward.   It will, essentially, be "business as usual" from Boulder from this point.

We are all excited for a successful Final Cutover this morning, and congratulate the many ADOA-ASET, IBM and State Agency Representatives, whose incredible collaboration and partnership were the catalyst that drove this project to this point.

The next posting will occur at 5:00 AM this morning.

 

 

 

Sunday, January 7, 2018 - 11:43

Good morning, 

At this time, the User Testing window for Dry Run #2 has expired, and the IBM team is working to bring the Phoenix system back up by no later than 1:00 PM this afternoon.

The Boulder system should also be back up and ready for user testing, utilizing test IP addresses, by 5:30 PM this afternoon.

We will have a follow up discussion on the events of this Dry Run during our Daily Agency User Testing Touch Point Call at 10:30 AM tomorrow.

Thank you to all of our agency partners who worked with us this morning on Dry Run #2.

This will be the last posting for this event.

 

Sunday, January 7, 2018 - 10:15

Good morning, 

The current time is 10:15 AM.   We will have another 30 minutes in our extended testing window for agencies, and testing appears to be, overall, successfully executed.

The testing window will conclude at 10:45 AM.

Thank you.

 

 

Sunday, January 7, 2018 - 09:07

Good morning, 

We are going to extend the testing window for one hour, and it will now end at 10:45 AM.

The next update will be provided at 10:00 AM.

Thank you.

Sunday, January 7, 2018 - 07:07

Good morning -

The current time is 7:07 AM

The latest update is that the IODEF file that was not saved to Boulder issue has been resolved and the LPARs are beginning to be brought up in Boulder.

Unfortunately, this snag has pushed our testing window back one hour, and the new approximate start time for agency testing is now 7:30 AM.

We still do not know at the present time whether we will be able to add some testing time to the back end of the schedule, but as soon as this is determined, you will be made aware.

We appreciate your patience, and apologize for the delay.

The next update will be posted at 8:00 AM. 

 

Sunday, January 7, 2018 - 06:02

Good morning -

The time is 6:02 AM.

Due to an issue with the Storage Team, we have to push back the agency testing window.  Currently, we are predicting a start time for agency testing of 6:50 AM.

Another update will be posted before 7:00 AM, if we get another update before then.

Apologies for the delay.

Pages